Find out inverse Laplace transforms of the following provided the

systems assuming a) system is stable, b) assuming system is causal.

2

H(s) -

(5-1)(3-3)

1

H(s) -

(s + 1)(3+3)

1

H(s) =

(s + 1)(-3)


1
Expert's answer
2021-06-10T03:46:56-0400

Given:

"r(t) \u2217 h(t) = b(t)"

"\implies R(s)H(s) = B(s)"

"R(s) =\frac{1}{s^2}"

"B(s) = \frac{s}{s^2+5s+6}"

"\implies H(s) =\frac{s^3}{s^2+5 s+6}"

"Y (s) = U(s)X(s)"

"\implies Y (s) =\frac{s^2}{s^2+5 s+6}"

Taking Inverse Laplace Transform

"\implies y(t) = 4e^{−2t}u(t) − 9e^{−3t}u(t) + δ(t)"